For example I want to calculate national government I have total income(Y) and consumption(C) and have investment(I) also I have import(IM ) but I don't have (G) and (EX)... EXANPLE : If y=1137179925 and c=68439 and I=22783 import=39252 find G? And export
1
Expert's answer
2017-10-06T13:26:06-0400
Y = C + I + G + EX – IM
1137179925 = 68439 + 22783 + G + EX – 39252
G + EX = 1137127955